### v1

@@ Abstract
We propose a new statistical ensemble of toric bases for elliptic Calabi-Yaus used in F-theory models, by focusing on only the convex hull of the base, i.e., the base polytope. This physically motivated coarse-graining greatly simplifies the combinatorial complexity of the part of the 4D F-theory landscape with toric bases. We develop a Monte Carlo approach that randomly samples the base polytopes within fixed boxes, with proper statistical weights. We first apply the algorithm to the set of 2d base polytopes, generating an enlarged set of toric 2d bases that include certain types of codimension-two (4,6) points, and we validate our approach against exact numbers. We then explore the set of 3d base polytopes which fit in a set of “maximal” 3d boxes, and estimate the total number of inequivalent 3d base polytopes to be $\sim 10^{85}$$10^{90}$. We provide statistical data such as the distribution of non-Higgsable gauge groups on these bases. Amusingly, a similar method can also be applied to generate reflexive polytopes in various dimensions. In both the reflexive and base polytope cases, the number of relevant polytopes obeys a Gaussian distribution as a function of the number of vertices, which can be understood in terms of other results on random polytopes in the math literature.
preprint: September 16, 2025   preprint: MIT-CTP/5924
@@ 1 Introduction
    On a fixed base $B_{d}$, classifying different elliptic fibration structures leading to topologically distinct elliptic Calabi-Yau $(d+1)$-fold $X_{d+1}$. Physically, these distinct choices can give rise to different enhanced (tuned) geometric gauge groups, matter curves and geometric Yukawa couplings for different $d$. While the Weierstrass model (1) can be directly tuned to realize a variety of nonabelian gauge factors, matter representations, and low-rank abelian gauge factors, the full set of possible fibrations is not completely understood; in the toric context, there has been some partial exploration of the tuning possibilities for CY3 hypersurfaces in reflexive polytopes Huang:2018esr ; Huang:2019pne ; Abbasi-nt . One can also potentially add frozen singularities (O7+ branes) Tachikawa:2015wka ; Bhardwaj:2018jgp ; Morrison:2023hqx ; Oehlmann:2024cyn .
@@
The first step in developing a comprehensive global picture of the 4d F-theory landscape is thus the classification of the set of 3d compact complex bases that support an elliptic Calabi-Yau fourfold. Unlike in the case of 2d bases, this set of bases is not completely classified (although it is known to be finite Di_Cerbo_2021 ), and even the relatively controlled domain of toric 3d bases is only partially understood. In this paper, we revisit the problem of classifying toric base threefolds $B_{3}$ for 4d F-theory. In fact, if one tries to classify all the possible bases up to isomorphism, when $h^{1,1}(B_{3})$ is large, there is a huge number of different bases related to each other via flips and flops. For example, the base $B_{3}$ of the elliptic CY4 with the largest $h^{1,1}=303148$ has at least $7.5\times 10^{45766}$ flips and flop phases Wang:2020gmi . Previous studies of the set of toric threefold bases have included this kind of redundancy.
@@
Of course, a given base polytope $\Delta_{b}$ may not lead to a smooth toric threefold. In other words, the base may possess terminal singularities for any triangulation. A simple example of this case would be the base polytope with vertices $\{(1,0,0)$, $(0,1,0)$, $(0,0,1)$, (1,1,2)}(-1,-1,-2)\}. In this paper we would generally allow such kind of base polytope to be a part of our enlarged base ensemble. Note that such singular bases are a somewhat generic feature of the 3d bases expected in the F-theory landscape.
@@
For the set of 3d boxes in which we sample the base polytopes, we generate a list of 4553 “maximal” boxes $\mathcal{B}$ that can be thought as the dual of “minimal” $\mathcal{G}$-polytopes (without proving their completeness). The minimal $\mathcal{G}$-polytopes can be thought as starting points of blow-ups in the minimal model of toric threefold bases, which can be smooth, and they are listed in the attached file “3d-G-polytopes.txt”. In this way, any base polytope $\Delta_{b}$ in the box $\mathcal{B}$ with origin contained in the interior of $\Delta_{b}$ corresponds to a valid compact F-theory base at finite distance in moduli space [^1] For the cases with infinite distance, see Grimm:2018ohb ; Lee:2021usk ; Alvarez-Garcia:2023gdd ; Alvarez-Garcia:2023qqj ; Chen:2024cvc for recent developments using emergent extra dimension or strings..
@@
Given that the total number of 3d maximal boxes is around $\sim 5\times 10^{3}$, it is reasonable to give an upper bound estimation of the total number of 3d base polytopes as $\sim 10^{90}$. and we compute the statistics of their Hodge numbers and the number of rigid (automatically present) gauge groups on each threefold base, including the ones with $E_{7}$ gauge groups that could potentially realize an $E_{7}$ GUT model following Li:2021eyn ; Li:2022aek ; Li:2023dya , as well as $E_{8}$ gauge groups that are the dominant rigid group containing the Standard Model gauge group as a subgroup. Interestingly, the total number of base polytopes in each of the four (rather different) boxes that we consider as example cases are all of the same order of magnitude, despite a large difference in the sizes of these boxes. We also find that the non-Higgsable $E_{7}$ gauge groups are more common in the Monte Carlo approaches in the boxes $\mathcal{B}_{6}(21,129)$, $\mathcal{B}_{6}(42,258)$ and $\mathcal{B}_{6}(63,387)$, for bases in the range of $h^{1,1}(B_{3})\leq 140000$, while the number of $E_{8}$ factors is much greater in the larger boxes.
To justify the statistical approach, we have also applied the Monte Carlo program to the case of 2d base polytopes, in which we can compare our results with a rigorous enumeration of such base polytopes. In the 2d cases, the set of minimal $\mathcal{G}$-polytopes are exactly $\mathbb{P}^{2}$, $\mathbb{F}_{0}$ and $\mathbb{F}_{n}$ $(12\geq n\geq 2)$, which are the starting points of blow-ups for 2d bases of F-theory Grassi:1991tjo ; Gross:1993fd ; Morrison:2012js . For 2d bases, the triangulation is unique, and from this perspective the set of allowed base polytopes would be larger than the set of 61539 2d bases in Morrison:2012js . In fact, this ensemble includes smooth bases with certain type of codimension-two (4,6) singularities in the elliptic fibration. We estimate the total number of 2d base polytopes to be $N_{\rm tot,2d}\approx 2.9\times 10^{9}$. We also provide various consistency checks in subsets, i.e. smaller 2d boxes defined relative to a $k$-dual set of minimal polytopes, for which the complete set of 2d base polytopes can be enumerated.
@@
In Section 4 we present the most important $d=3$ case. We first generate a (partial) list of minimal $\mathcal{G}$-polytopes that are dual to the maximal 3d boxes used later, including $\mathcal{G}$-polytopes with vertices $\{(1,0,0),(0,1,0),(0,0,1),(-1,-p,-q)\}$ and the more general cases, in Section 4.1. In Section 4.2 we present and analyze the Monte Carlo data sampled in the largest box $\mathcal{B}_{6}(84,516)$, and then do the same for three other boxes in Section 4.3. We plot the average number of each kind of non-Higgsable gauge groups as a function of $h^{1,1}(B_{3})$ in Section 4.4, and then discuss the geometric and physical meaning of the ensemble of 3d bases we get in Section 4.5.
@@ 2.1 Sampling polytopes in a box
The general combinatoric questions we would like to address are: what is the total number of possible $d$-dimensional base polytopes associated with compact toric $d$-folds that can support an elliptic fibration, how can we randomly sample them, and what are their statistics. In this section we take the dimension $d$ to be general.
@@
The first and most significant redundancy factor, denoted by $R(\Delta_{b},m,n)$, describes the total number of the ways to choose $n$ points such that their convex hull forms the same polytope $\Delta_{b}$. When $m<n$, (nm)(n-m) points out of $(v_{1},v_{2},\dots v_{n})$ are general primitive points located inside the polytope $\Delta_{b}$, including the boundary (and vertices) of $\Delta_{b}$. In this case, there are additional redundancy factors due to the fact that these (nm)(n-m) vertices can be arbitrarily chosen inside $\Delta_{b}$, and these different choices all give rise to the same $\Delta_{b}$. We define $l(\Delta_{b})$ to be the number of points inside $\Delta_{b}$, including the boundary of $\Delta_{b}$.
@@
A rough heuristic explanation for this formula can be understood as follows: when $l(\Delta_{b})\gg n$, the probability that one point is chosen twice or more is suppressed by the factor $n/l(\Delta_{b})$. Therefore, in a good approximation, we can omit such contributions, and assume that $n$ choices result in $n$ different points. In such circumstance, there are $m$ points chosen to be the vertices and the remaining points lie in the other $l(\Delta_{b})-m$ locations, and it is straightforward to check that $R=(l(\Delta_{b})-m)^{n-m}m!\binom{n}{m}$.
The second independent redundancy factor comes from the possible $GL(d,\mathbb{Z})$ transformations $v_{i}^{\prime}=M\cdot v_{i}$ $(i=1,\dots,n)$, such that the transformed vertices $v_{i}^{\prime}$ $(i=1,\dots,n)$ still lie inside $\mathcal{B}$, and $v_{i}^{\prime}$ is not a permutation of the old set of vertices $v_{i}$ (if it is a permutation, the redundancy is already taken into account in $R$). We denote the number of such $GL(d,\mathbb{Z})$ transformations modding out the $S_{n}$ permutations as $R_{GL(d,\mathbb{Z})}$. The estimated total number of $\Delta_{b}$ should be further divided by $1+R_{GL(d,\mathbb{Z})}$, due to the existence of $1+R_{GL(d,\mathbb{Z})}$ equivalent bases with different shapes in $\mathcal{B}$. After these two redundancy factors are taken into account, the estimated total number of bases in the Monte Carlo approach is
@@
and the total number of bases
@@ 2.2 Set of maximal boxes
Imagine one starts from a very large box; a random polytope inside the box would then likely fail to satisfy the $\mathcal{G}$-polytope condition, as the $\mathcal{G}$ polytope is likely to be very small and not contain origin $\vec{0}$ in its interior. To increase the efficiency, in the paper we use an alternative approach, such that we first try to classify the different minimal $\mathcal{G}$ polytopes for compact $d$-dimensional toric bases. Then for each of these minimal $\mathcal{G}$ polytopes, we define a box $\mathcal{B}_{6}(\mathcal{G})$ that is defined as the six times polar dual of $\mathcal{G}$; more generally, we define
@@
In fact, the “minimal condition” in the condition 3 above exactly coincides with the criteria for a minimal model in the toric minimal model program, and $\mathcal{G}$ can be equivalently thought as a toric variety that cannot be blown down. Due to the condition 1, this toric variety should also be a valid, possibly singular base of an elliptic Calabi-Yau $(d+1)$-fold.
@@ 2.3 Gaussian distributions
after we absorb all the constant terms.
One important variable, the number of vertices $m$, has a large effect on $\ln(R(l(\Delta_{b}),m,n))$. As can be seen from examples, this variable $m$ also approximately fits into a Gaussian distribution333nn and $l(\Delta_{b})$ should have effects as well, but the $n$ dependence should primarily affect the regime sampled, and not the net weights, and the range of variation in $l(\Delta_{b})/R\sim 1$ is not very wide either. Hence we primarily focus on the dependence on the single variable that is the number of vertices $m$.. More rigorously, it has been proven in Theorem 1.2 of vu2006central that for an ensemble of random convex polytopes in a convex set in $\mathbb{R}^{d}$, quantities such as the number of vertices obey a form of the central limit theorem and give a normal distribution. We will see in Section 3 and Section 4 two different Gaussian distributions that can be anticipated following this argument. First, the numbers of samplings in a Monte Carlo run obey a Gaussian distribution with respect to the number of vertices $m$. Second, the numbers of samplings obey a Gaussian distribution with respect to the $\log(w_{i})$. The first of these results can be seen as a discrete analogue of the results in vu2006central , while the second can be understood from (24), with the variation largely dominated by terms linear in $m$.
@@ 3.1 Set of all 2d base polytopes
We have implemented both of these algorithms in Julia and run them on a computing cluster for all values of $k\leq 6$. For $k\leq 4$, both algorithms run to completion, and we have checked that the results agree. Since the top-down algorithm is run separately for each box, while the bottom-up algorithm automatically includes all boxes, this agreement is a strong check on the methodology. For $k=5$ the top-down algorithm also ran to completion; the total number of polytopes for each value of $k\leq 5$ is given in Table 2. For $k=6$, the numbers of polytopes become quite large ($10^{9}$), and we have only run the computation from the bottom up to $n\sim 30$ and from the top down to $n\sim 130$. These results for $k=6$ and $k=5$ are described in more detail in the following subsections in comparison to the results of the Monte Carlo analysis.

### v2

@@ Abstract
We propose a new statistical ensemble of toric bases for elliptic Calabi-Yaus used in F-theory models, by focusing on only the convex hull of the base, i.e., the base polytope. This physically motivated coarse-graining greatly simplifies the combinatorial complexity of the part of the 4d F-theory landscape with toric bases. We develop a Monte Carlo approach that randomly samples the base polytopes within fixed boxes, with proper statistical weights. We first apply the algorithm to the set of 2d base polytopes, generating an enlarged set of toric 2d bases that include certain types of codimension-two (4,6) points, and we validate our approach against exact numbers. We then explore the set of 3d base polytopes which fit in a set of “maximal” 3d boxes, and estimate the total number of inequivalent 3d base polytopes to be $10^{85}$$10^{90}$. We provide statistical data such as the distribution of non-Higgsable gauge groups on these bases. Amusingly, a similar method can also be applied to generate reflexive polytopes in various dimensions. In both the reflexive and base polytope cases, the number of relevant polytopes obeys a Gaussian distribution as a function of the number of vertices, which can be understood in terms of other results on random polytopes in the math literature.
preprint: December 6, 2025   preprint: MIT-CTP/5924
@@ 1 Introduction
    On a fixed base $B_{d}$, classifying different elliptic fibration structures leading to topologically distinct elliptic Calabi-Yau $(d+1)$-folds $X_{d+1}$. Physically, these distinct choices can give rise to different enhanced (tuned) geometric gauge groups, matter curves and geometric Yukawa couplings for different $d$. While the Weierstrass model (1) can be directly tuned to realize a variety of nonabelian gauge factors, matter representations, and low-rank abelian gauge factors, the full set of possible fibrations is not completely understood; in the toric context, there has been some partial exploration of the tuning possibilities for CY3 hypersurfaces in reflexive polytopes Huang:2018esr; Huang:2019pne; Abbasi-nt. One can also potentially add frozen singularities (O7+ branes) Tachikawa:2015wka; Bhardwaj:2018jgp; Morrison:2023hqx; Oehlmann:2024cyn.
@@
The first step in developing a comprehensive global picture of the 4d F-theory landscape is thus the classification of the set of 3d compact complex bases that support an elliptic Calabi-Yau fourfold. Unlike in the case of 2d bases, this set of bases is not completely classified (although it is known to be finite Di_Cerbo_2021), and even the relatively controlled domain of toric 3d bases is only partially understood. In this paper, we revisit the problem of classifying toric base threefolds $B_{3}$ for 4d F-theory. In fact, if one tries to classify all the possible bases up to isomorphism, when $h^{1,1}(B_{3})$ is large, there is a huge number of different bases related to each other via flips and flops. For example, the base $B_{3}$ of the elliptic CY4 with the largest $h^{1,1}=303148$ has at least $7.5\times 10^{45766}$ flip and flop phases Wang:2020gmi. Previous studies of the set of toric threefold bases have included this kind of redundancy.
@@
Of course, a given base polytope $\Delta_{b}$ may not lead to a smooth toric threefold. In other words, the base may possess terminal singularities for any triangulation. A simple example of this case would be the base polytope with vertices $\{(1,0,0)$, $(0,1,0)$, $(0,0,1)$, (1,1,2)}(-1,-1,-2)\}. In this paper we would generally allow such kinds of base polytope to be a part of our enlarged base ensemble. Note that such singular bases are a somewhat generic feature of the 3d bases expected in the F-theory landscape.
@@
For the set of 3d boxes in which we sample the base polytopes, we generate a list of 4553 “maximal” boxes $\mathcal{B}$ that can be thought as the dual of “minimal” $\mathcal{G}$-polytopes (without proving their completeness), defined as
$\mathcal{B}=\{v|\langle u,v\rangle\geq-6\ ,\ u\in\mathcal{G}\}\,.$ (4)
The minimal $\mathcal{G}$-polytopes can be thought as starting points of blow-ups in the minimal model of toric threefold bases, which can be smooth [^1] The list of these polytopes can be found on https://doi.org/10.5281/zenodo.17761959 in 3d/3d-G-polytopes.txt.. In this way, any base polytope $\Delta_{b}$ in the box $\mathcal{B}$ with origin contained in the interior of $\Delta_{b}$ corresponds to a valid compact F-theory base at finite distance in moduli space [^2] For the cases with infinite distance, see Grimm:2018ohb; Lee:2021usk; Alvarez-Garcia:2023gdd; Alvarez-Garcia:2023qqj; Chen:2024cvc for recent developments using emergent extra dimensions or strings.. A similar construction of “maximal” polytopes was used in Avram:1997rs to classify reflexive polytopes.
@@
Given that the total number of 3d maximal boxes is around $\sim 5\times 10^{3}$, it is reasonable to give an upper bound estimation of the total number of 3d base polytopes as $\sim 10^{90}$. We compute the statistics of the Hodge numbers and the number of rigid (automatically present) gauge groups on threefold bases in these boxes, including the ones with $E_{7}$ gauge groups that could potentially realize an $E_{7}$ GUT model following Li:2021eyn; Li:2022aek; Li:2023dya, as well as $E_{8}$ gauge groups that are the dominant rigid group containing the Standard Model gauge group as a subgroup. Interestingly, the total number of base polytopes in each of the four (rather different) boxes that we consider as example cases are all of the same order of magnitude, despite a large difference in the sizes of these boxes. We also find that the non-Higgsable $E_{7}$ gauge groups are more common in the Monte Carlo approaches in the boxes $\mathcal{B}_{6}(21,129)$, $\mathcal{B}_{6}(42,258)$ and $\mathcal{B}_{6}(63,387)$, for bases in the range of $h^{1,1}(B_{3})\leq 140000$, while the number of $E_{8}$ factors is much greater in the larger boxes.
To confirm the validity of the statistical approach, we have also applied the Monte Carlo program to the case of 2d base polytopes, in which we can compare our results with a rigorous enumeration of such base polytopes. In the 2d cases, the set of minimal $\mathcal{G}$-polytopes are exactly $\mathbb{P}^{2}$, $\mathbb{F}_{0}$ and $\mathbb{F}_{n}$ $(12\geq n\geq 2)$, which are the starting points of blow-ups for 2d bases of F-theory Grassi:1991tjo; Gross:1993fd; Morrison:2012js. For 2d bases, the triangulation is unique, and from this perspective, the set of allowed toric base polytopes is larger than the set of 61539 2d bases in Morrison:2012js. In fact, this ensemble includes smooth bases with certain type of codimension-two (4,6) singularities in the elliptic fibration. We estimate the total number of 2d base polytopes to be $N_{\rm tot,2d}\approx 2.9\times 10^{9}$. We also provide various consistency checks in subsets, i.e. smaller 2d boxes defined relative to a $k$-dual set of minimal polytopes, for which the complete set of 2d base polytopes can be enumerated.
@@
In Section 4 we present the most important $d=3$ case. We first generate a (partial) list of minimal $\mathcal{G}$-polytopes that are dual to the maximal 3d boxes used later, including $\mathcal{G}$-polytopes with vertices $\{(1,0,0),(0,1,0),(0,0,1),(-1,-p,-q)\}$ and the more general cases, in Section 4.1. In Section 4.2 we present and analyze the Monte Carlo data sampled in the largest box $\mathcal{B}_{6}(84,516)$, and then do the same for three other boxes in Section 4.3. We plot the average number of each kind of non-Higgsable gauge group factors as a function of $h^{1,1}(B_{3})$ in Section 4.4, and then discuss the geometric and physical meaning of the ensemble of 3d bases we get in Section 4.5.
@@
Various software programs and associated data files used to produce some of the results in this paper are included in a Zenodo archive https://doi.org/10.5281/zenodo.17761959.
@@ 2.1 Sampling polytopes in a box
The general combinatoric questions we would like to address are: “What is the total number of possible $d$-dimensional base polytopes associated with compact toric $d$-folds that can support an elliptic fibration?,” “How can we randomly sample them?,” and “What are their statistics?”. In this section we take the dimension $d$ to be general.
@@
The first and most significant redundancy factor, denoted by $R(\Delta_{b},m,n)$, describes the total number of ways to choose $n$ points such that their convex hull forms the same polytope $\Delta_{b}$ with $m$ vertices. When $m<n$, (nm)(n-m) points out of $(v_{1},v_{2},\dots v_{n})$ are general primitive points located inside the polytope $\Delta_{b}$, including the boundary (and vertices) of $\Delta_{b}$. In this case, there are additional redundancy factors due to the fact that these (nm)(n-m) points can be arbitrarily chosen inside $\Delta_{b}$, and these different choices all give rise to the same $\Delta_{b}$. We define $l(\Delta_{b})$ to be the number of points inside $\Delta_{b}$, including the boundary of $\Delta_{b}$.
@@
A rough heuristic explanation for this formula can be understood as follows: when $l(\Delta_{b})\gg n$, the probability that one point is chosen twice or more is suppressed by the factor $n/l(\Delta_{b})$. Therefore, in a good approximation, we can omit such contributions, and assume that $n$ choices result in $n$ different points. In such circumstances, there are $m$ points chosen to be the vertices and the remaining points lie in the other $l(\Delta_{b})-m$ locations, and it is straightforward to check that $R=(l(\Delta_{b})-m)^{n-m}m!\binom{n}{m}$.
The second independent redundancy factor comes from the possible $GL(d,\mathbb{Z})$ transformations $v_{i}^{\prime}=M\cdot v_{i}$ $(i=1,\dots,n)$, such that the transformed vertices $v_{i}^{\prime}$ $(i=1,\dots,n)$ still lie inside $\mathcal{B}$, and $v_{i}^{\prime}$ is not a permutation of the old set of vertices $v_{i}$ (if it is a permutation, the redundancy is already taken into account in $R$). We denote the number of such $GL(d,\mathbb{Z})$ transformations, modding out the $S_{n}$ permutations, as $R_{GL(d,\mathbb{Z})}$. The estimated total number of $\Delta_{b}$ should be further divided by $1+R_{GL(d,\mathbb{Z})}$, due to the existence of $1+R_{GL(d,\mathbb{Z})}$ equivalent bases with different shapes in $\mathcal{B}$. After these two redundancy factors are taken into account, the estimated total number of bases in the Monte Carlo approach is
@@
and the total number of bases is then
@@ 2.2 Set of maximal boxes
Imagine one starts from a very large box; a random polytope inside the box would then likely fail to satisfy the $\mathcal{G}$-polytope condition, as the $\mathcal{G}$ polytope is likely to be very small and not contain the origin $\vec{0}$ in its interior. To increase efficiency, in the paper we use an alternative approach, such that we first try to classify the different minimal $\mathcal{G}$ polytopes for compact $d$-dimensional toric bases. Then for each of these minimal $\mathcal{G}$ polytopes, we define a box $\mathcal{B}_{6}(\mathcal{G})$ that is defined as the six times polar dual of $\mathcal{G}$; more generally, we define
@@
In fact, the “minimal condition” in the condition 3 above exactly coincides with the criteria for a minimal model in the toric minimal model program, and $\mathcal{G}$ can be equivalently thought of as a toric variety that cannot be blown down. Due to the condition 1, this toric variety should also be a valid, possibly singular base of an elliptic Calabi-Yau $(d+1)$-fold.
@@ 2.3 Gaussian distributions
after we absorb all the constant terms. One can also convert the formula into the $\log_{10}$ base.
One important variable, the number of vertices $m$, has a large effect on $\ln(R(l(\Delta_{b}),m,n))$. As can be seen from examples, this variable $m$ also approximately fits into a Gaussian distribution444nn and $l(\Delta_{b})$ should have effects as well, but the $n$ dependence should primarily affect the regime sampled, and not the net weights, and the range of variation in $l(\Delta_{b})/R\sim 1$ is not very wide either. Hence we primarily focus on the dependence on the single variable that is the number of vertices $m$.. More rigorously, it has been proven in Theorem 1.2 of vu2006central that for an ensemble of random convex polytopes in a convex set in $\mathbb{R}^{d}$, quantities such as the number of vertices obey a form of the central limit theorem and give a normal distribution. We will see in Section 3 and Section 4 two different Gaussian distributions that can be anticipated following this argument. First, the numbers of samplings in a Monte Carlo run obey a Gaussian distribution with respect to the number of vertices $m$. Second, the numbers of samplings obey a Gaussian distribution with respect to the $\log(w_{i})$. The first of these results can be seen as a discrete analogue of the results in vu2006central, while the second can be understood from (25), with the variation largely dominated by terms linear in $m$.
@@ 3.1 Set of all 2d base polytopes
We have implemented both of these algorithms in Julia and run them on a computing cluster for all values of $k\leq 6$ [^5] See the files on https://doi.org/10.5281/zenodo.17761959 in the directory “2d-code”.. For $k\leq 4$, both algorithms run to completion, and we have checked that the results agree. Since the top-down algorithm is run separately for each box, while the bottom-up algorithm automatically includes all boxes, this agreement is a strong check on the methodology. For $k=5$ the top-down algorithm also ran to completion; the total number of polytopes for each value of $k\leq 5$ is given in Table 2. For $k=6$, the numbers of polytopes become quite large ($10^{9}$), and we have only run the computation from the bottom up to $n\sim 30$ and from the top down to $n\sim 130$. These results for $k=6$ and $k=5$ are described in more detail in the following subsections in comparison to the results of the Monte Carlo analysis.